OpenStudy (anonymous):

compute: 4 x 2^2 -10 / 2 +5

OpenStudy (anonymous):

\[4 * 2^{2} -10 \div2+5\]

OpenStudy (blurbendy):

Parentheses Exponents Multiplication Division Addition Subtraction PEMDAS Start with the Exponents. 2^2 = 4 4 * 4 - 10 / 2 + 5 4 * 4 = 16 16 - 10 / 2 + 5 16 - 5 + 5 11 + 5 = 16 Did you follow?

OpenStudy (anonymous):

Yes i did

OpenStudy (blurbendy):

Yay!

OpenStudy (anonymous):

but on another ? i followed the rules and came up wrong ..... 3+5(10-2) - 8(2+3) I got 12 as an answer & it was wrong